# Truth or Twist? Optimal Model Selection for Reliable Label Flipping Evaluation in LLM-based Counterfactuals

### v2

@@ Abstract
Counterfactual examples are widely employed to enhance the performance and robustness of large language models (LLMs) through counterfactual data augmentation (CDA). However, the selection of the judge model used to evaluate label flipping, the primary metric for assessing the validity of generated counterfactuals for CDA, yields inconsistent results. To decipher this, we define four types of relationships between the counterfactual generator and judge models: being the same model, belonging to the same model family, being independent models, and having an distillation relationship. Through extensive experiments involving two state-of-the-art LLM-based methods, three datasets, four generator models, and 15 judge models, complemented by a user study $(n=90)$, we demonstrate that judge models with an independent, non-fine-tuned relationship to the generator model provide the most reliable label flipping evaluations. [^1] Code and evaluation results are available at: https://github.com/qiaw99/failure Relationships between the generator and judge models, which are closely aligned with the user study for CDA, result in better model performance and robustness. Nevertheless, we find that the gap between the most effective judge models and the results obtained from the user study remains considerably large. This suggests that a fully automated pipeline for CDA may be inadequate and requires human intervention.
@@ 1 Introduction
Counterfactual examples are minimally altered versions of original inputs that flip the initial label Miller (2019); Ross et al. (2021); Madsen et al. (2022). They serve as a valuable approach for CDA aimed at improving model robustness and performance Liu et al. (2021); Dixit et al. (2022); Balashankar et al. (2023); Agrawal et al. (2025). We want to emphasize the subtle yet significant distinction between counterfactuals used for explaining model predictions and those used for CDA. In the former, the objective is to flip the model’s prediction, while the goal of CDA is to flips the ground truth label (Figure 6 in Appendix B). [^2] In this study, we mainly focus on the latter type of counterfactuals used for CDA. To evaluate the effectiveness and validity of the LLM-generated counterfactuals for CDA, the label flip rate (LFR) is a common metric of choice Ge et al. (2021). It is the percentage of valid counterfactuals where the ground truth labels are flipped out of the total number of instances. LFR of counterfactuals can be evaluated using either the same model that generates the counterfactual Bhattacharjee et al. (2024a, b); Wang et al. (2025a) or independent models Dixit et al. (2022); Balashankar et al. (2023). The optimal strategy for selecting models to evaluate the ground-truth validity of counterfactuals remains uncertain (Figure 1). This uncertainty, in turn, hampers efforts to enhance model robustness and performance through CDA, as noisy or erroneous labels may degrade model performance.
@@
In this work, we first define four types of relationships between the counterfactual generator model and the judge model: being the same model, independent models with and without fine-tuning on the target dataset, distilled models, and models from the same family (Figure 1). Secondly, we conduct comprehensive experiments to predict labels for counterfactuals generated by two state-of-the-art approaches across three datasets and four generator models, with 15 judge models. Thirdly, we undertake a user study to assess the validity of the generated counterfactuals in acquiring ground-truth LFR. Our finding shows that a judge model with an independent, non-fine-tuned relationship to the generator captures label flipping most effectively. Relationships between the generator and judge models that are most aligned with the user study, lead to improved model performance and robustness. Additionally, there remains a considerable gap between the performance of the best judge models and the results observed in the user study.
@@ 2 Background and Related Work
Initially, LLMs generate counterfactuals using established counterfactual generation approaches (Figure 4 in Appendix A). These generated counterfactuals should subsequently be validated—either by human annotators or judge models—as invalid counterfactuals bearing incorrect labels may degrade model performance Song et al. (2023). In practice, relying solely on human evaluation is both costly and inefficient; therefore, LLMs are commonly employed to validate the generated counterfactuals. Finally, valid counterfactuals are utilized as additional training data to enhance model performance and robustness Yang et al. (2021); Dixit et al. (2022).
@@
The literature identifies two primary ways to verify label flipping in edited inputs: (1) employing an independent model, distinct from the one producing counterfactuals; (2) utilizing the same LLM that generates counterfactuals, guided by carefully constructed prompts. Prior to, and even during the widespread adoption of LLMs, encoder-only models, e.g., DeBERTa Dixit et al. (2022), RoBERTa Ross et al. (2021); Balashankar et al. (2023); Treviso et al. (2023) or BERT Kaushik et al. (2020); Fern and Pope (2021); Robeer et al. (2021), were predominantly used to verify label flipping. This preference stems from their advanced performance in text classification tasks. In more recent work, the same LLMs are increasingly employed both to generate counterfactuals Bhattacharjee et al. (2024a, b); Wang et al. (2025a); Dehghanighobadi et al. (2025) and to evaluate them for label flipping, since their classification accuracy rivals that of fine-tuned encoder-only models.
@@ 3.2 Relationships
  • Distilled models ($\mathcal{R}_{dm}$): $LLM_{\mathcal{G}}$ and $LLM_{\mathcal{J}}$ have equal number of parameters and the same architecture. $LLM_{\mathcal{J}}$ is distilled and fine-tuned using synthetic data from a third model $LLM_{\mathcal{D}}$, which is more powerful and not part of the model family $\mathcal{F(M)}$ of generator and judge model.
@@ 4.1 Counterfactual Methods Selection
We select two state-of-the-art approaches based on LLMs that are shown to generate counterfactual examples efficiently and effectively: FIZLE Bhattacharjee et al. (2024a) and FLARE Bhattacharjee et al. (2024b). FIZLE first enables LLMs to identify key words within the input and then leverages these words to guide the generation of counterfactual examples. Meanwhile, FLARE generates counterfactuals by prompting LLMs in three steps: extracting latent features, identifying relevant words linked to those features, and modifying these words to produce counterfactual examples.
@@ 4.4 User Study
We recruit 90 native English speakers and, for each of the three datasets, randomly sample 45 indices. For each subset, i.e., generator-dataset pair (Table 3 in Appendix G), the counterfactuals generated by the corresponding generator model $LLM_{\mathcal{G}}$ for the selected indices are evaluated by two human annotators. If no majority label emerges from the labels provided by human annotators, we determine one of the two labels chosen by the two annotators, ensuring the ground-truth label is agreed upon by two people. [^5] The annotation guidelines and annotator information are provided in Appendix E. We further conduct an automatic evaluation in Appendix G.3 on the selected 45 counterfactuals as a sanity check to validate their representativeness of the overall distribution. Each annotator is given 15 counterfactuals, along with the set of possible labels given by the dataset, and tasked with selecting the optimal label. We report an inter-annotator agreement of Cohen’s $\kappa=0.55$. [^6] Although we employ two state-of-the-art methods – FIZLE and FLARE – to generate counterfactuals, they are not consistently perfect (at times failing to fully shift the semantics from the original to the target label) and sometimes produce ambiguous cases (Figure 12). This has an effect on the IAA being moderate which would likely improve with the development and use of more advanced counterfactual-generation methods. Lastly, we calculate the ground-truth LFR as the proportion of valid counterfactuals relative to the total number of instances (Table 4 in Appendix G).
@@ 4.5.2 Human Alignment Evaluation
To obtain the rankings, we first calculate LFR for human annotators and for each judge-generator model pair on each set of counterfactuals generated by given generator models, as reported in Table 4 of Appendix G. Next, we compute the difference ($\Delta$) between the human LFR and the LFR of each pair. A smaller difference indicates a better ranking (lower ranking value). This process results in a ranking for each judge-generator pair. Since these pairs correspond to specific relationships $\mathcal{R}(\mathcal{G},\mathcal{J})$, we average the rankings of all pairs sharing the same relationship to obtain the average ranking per relationship, as shown in Table 3 of Appendix G. Finally, we average these rankings across each generator model to produce the overall average rankings presented in Figure 2.
@@ 5 Results
As shown in Figure 3(a) and Figure 3(b), $LLM_{\mathcal{J}}$ with $\mathcal{R}_{imwo}$ relationship achieves the highest alignment with human annotators ($rank=4.15$, $r_{m/l}=3.5$, $\rho=0.47$). In contrast, $\mathcal{R}_{dm}$ ($rank=5.58$, $r_{m/l}=1$, $\rho=0.38$) or $\mathcal{R}_{imw}$ ($rank=5.86$, $r_{m/l}=0.23$, $\rho=0.29$) demonstrate poor alignment with human judgments. This may attribute to data contamination Li et al. (2025), as these models are either fine-tuned on the target dataset or share architectural similarities with $LLM_{\mathcal{G}}$. Such overlap could bias $LLM_{\mathcal{J}}$ in its evaluation of label flipping, potentially leading to either overestimating or underestimating the LFR. Notably, ensembling results from all judge models does not necessarily lead to better alignment, as it partially relies on results from suboptimal judge models. Additionally, we find the first two measures (average ranking and $r_{m/\ell}$) are moderately correlated, with a Spearman correlation coefficient of 0.67, indicating that $r_{m/\ell}$ also serves as a reliable metrics to capture alignment between model pairs and human judgments.
@@
We investigate how the choice of relationship affects CDA outcomes (Appendix H). We notice that when the LFR of $LLM_{\mathcal{J}}$, given its relationship to $LLM_{\mathcal{G}}$, aligns more closely with user study outcomes, the labels identified by $LLM_{\mathcal{J}}$ are associated with improved performance and greater robustness on both unseen and out-of-distribution data Kaushik et al. (2020); Gardner et al. (2020) (Appendix H). This association is particularly evident when these identified labels are treated as the ground-truth labels for counterfactuals, which subsequently serve as data points for CDA. This can be attributed to the fact that noisy and incorrect labels provided by judge models with suboptimal relationships contribute to performance deterioration Zhu et al. (2022); Song et al. (2023).
@@ 6 Conclusion
In this work, we emphasize the importance of the relationship between the counterfactual generator model and label flipping judge model in achieving LFR that align more closely with human annotations and in improving CDA outcomes. We further demonstrate that high downstream performance does not necessarily imply an effective judge model. Through extensive experiments, we identify that label flipping remains highly challenging. Additionally, the gap between the optimal relationship and the user study is considerably large, which indicates full automation of CDA falls short and human intervention should be considered.
@@ G.3 A Sanity Check
In our user study, we randomly select 45 counterfactuals from each dataset, generated individually by five counterfactual generator models $LLM_{\mathcal{G}}$4.4). These samples are then provided to human annotators to evaluate the effectiveness of the counterfactual, known as the simulatability test Wiegreffe et al. (2021); Feldhus et al. (2023). To perform a sanity check and validate the representativeness of these subsets, we conduct an additional automatic evaluation on the subset of input data from which the 45 counterfactuals were generated.

### v3

@@ Abstract
Counterfactual examples are widely employed to enhance the performance and robustness of large language models (LLMs) through counterfactual data augmentation (CDA). However, the selection of the judge model used to evaluate label flipping, the primary metric for assessing the validity of generated counterfactuals for CDA, yields inconsistent results. To decipher this, we define four types of relationships between the counterfactual generator and judge models: being the same model, belonging to the same model family, being independent models, and having an distillation relationship. Through extensive experiments involving two state-of-the-art LLM-based methods, three datasets, four generator models, and 15 judge models, complemented by a user study $(n=90)$, we demonstrate that judge models with an independent, non-fine-tuned relationship to the generator model provide the most reliable label flipping evaluations. [^1] Code and evaluation results are available at: https://github.com/qiaw99/truth-or-twist Relationships between the generator and judge models, which are closely aligned with the user study for CDA, result in better model performance and robustness. Nevertheless, we find that the gap between the most effective judge models and the results obtained from the user study remains considerably large. This suggests that a fully automated pipeline for CDA may be inadequate and requires human intervention.
@@ 1 Introduction
Counterfactual examples are minimally altered versions of original inputs that flip the initial label Miller (2019); Ross et al. (2021); Madsen et al. (2022). They serve as a valuable approach for CDA aimed at improving model robustness and performance Liu et al. (2021); Dixit et al. (2022); Balashankar et al. (2023); Agrawal et al. (2025). We want to emphasize the subtle yet significant distinction between counterfactuals used for explaining model predictions and those used for CDA. In the former, the objective is to flip the model’s prediction, while the goal of CDA is to flip the ground truth label (Figure 6 in Appendix B). [^2] In this study, we mainly focus on the latter type of counterfactuals used for CDA. To evaluate the effectiveness and validity of the LLM-generated counterfactuals for CDA, the label flip rate (LFR) is a common metric of choice Ge et al. (2021). It is the percentage of valid counterfactuals where the ground truth labels are flipped out of the total number of instances. LFR of counterfactuals can be evaluated using either the same model that generates the counterfactual Bhattacharjee et al. (2024a, b); Wang et al. (2025a) or independent models Dixit et al. (2022); Balashankar et al. (2023). The optimal strategy for selecting models to evaluate the ground-truth validity of counterfactuals remains uncertain (Figure 1). This uncertainty, in turn, hampers efforts to enhance model robustness and performance through CDA, as noisy or erroneous labels may degrade model performance.
@@
In this work, we first define four types of relationships between the counterfactual generator model and the judge model: being the same model, independent models with and without fine-tuning on the target dataset, distilled models, and models from the same family (Figure 1). Secondly, we conduct comprehensive experiments to predict labels for counterfactuals generated by two state-of-the-art approaches across three datasets and four generator models, with 15 judge models. Thirdly, we undertake a user study to assess the validity of the generated counterfactuals in acquiring a ground-truth LFR.
We find that a judge model with an independent, non-fine-tuned relationship to the generator captures label flipping most effectively. Relationships between the generator and judge models that are most aligned with the user study, lead to improved model performance and robustness. Additionally, there remains a considerable gap between the performance of the best judge models and the results observed in the user study.
@@ 2 Background and Related Work
At the start of a CDA process, LLMs generate counterfactuals using established counterfactual generation approaches (Figure 4 in Appendix A). These generated counterfactuals should subsequently be validated – either by human annotators or judge models – as invalid counterfactuals bearing incorrect labels may degrade model performance Song et al. (2023). In practice, relying solely on human evaluation is both costly and inefficient; therefore, LLMs are commonly employed to validate the generated counterfactuals. Finally, valid counterfactuals are utilized as additional training data to enhance model performance and robustness Yang et al. (2021); Dixit et al. (2022).
@@
The literature identifies two primary ways to verify label flipping in edited inputs: (1) employing an independent model, distinct from the one producing counterfactuals; (2) utilizing the same LLM that generates counterfactuals, guided by carefully constructed prompts. Prior to, and even during the widespread adoption of LLMs, encoder-only models, e.g., DeBERTa Dixit et al. (2022), RoBERTa Ross et al. (2021); Balashankar et al. (2023); Treviso et al. (2023) or BERT Kaushik et al. (2020); Fern and Pope (2021); Robeer et al. (2021), were predominantly used to verify label flipping. This preference stems from their superior performance in text classification tasks. In more recent work, the same LLMs are increasingly employed both to generate counterfactuals Bhattacharjee et al. (2024a, b); Wang et al. (2025a); Dehghanighobadi et al. (2025) and to evaluate them for label flipping, since their classification accuracy rivals that of fine-tuned encoder-only models.
@@ 3.2 Relationships
  • Distilled models ($\mathcal{R}_{dm}$): $LLM_{\mathcal{G}}$ and $LLM_{\mathcal{J}}$ have an equal number of parameters and the same architecture. $LLM_{\mathcal{J}}$ is distilled and fine-tuned using synthetic data from a third model $LLM_{\mathcal{D}}$, which is more powerful and not part of the model family $\mathcal{F(M)}$ of the generator and judge model.
@@ 4.1 Counterfactual Methods Selection
We select two state-of-the-art approaches based on LLMs that are shown to generate counterfactual examples efficiently and effectively: FIZLE Bhattacharjee et al. (2024a) and FLARE Bhattacharjee et al. (2024b). FIZLE first prompts LLMs to identify key words within the input and then leverages these words to guide the generation of counterfactual examples. Meanwhile, FLARE generates counterfactuals by prompting LLMs in three steps: extracting latent features, identifying relevant words linked to those features, and modifying these words to produce counterfactual examples.
@@ 4.4 User Study
We recruit 90 native English speakers and, for each of the three datasets, randomly sample 45 indices. For each subset, i.e., a generator-dataset pair (Table 3 in Appendix G), the counterfactuals generated by the corresponding generator model $LLM_{\mathcal{G}}$ for the selected indices are evaluated by two human annotators. If no majority label emerges from the labels provided by human annotators, we break the tie ourselves by selecting one of the two annotated labels, ensuring the ground-truth label is agreed upon by two people. [^5] The annotation guidelines and annotator information are provided in Appendix E. We further conduct an automatic evaluation in Appendix G.3 on the selected 45 counterfactuals as a sanity check to validate their representativeness of the overall distribution. Each annotator is given 15 counterfactuals, along with the set of possible labels given by the dataset, and tasked with selecting the optimal label. We report an inter-annotator agreement of Cohen’s $\kappa=0.55$. [^6] Although we employ two state-of-the-art methods – FIZLE and FLARE – to generate counterfactuals, they are not consistently perfect (at times failing to fully shift the semantics from the original to the target label) and sometimes produce ambiguous cases (Figure 12). This has an effect on the IAA being moderate which would likely improve with the development and use of more advanced counterfactual generation methods. Lastly, we calculate the ground-truth LFR as the proportion of valid counterfactuals relative to the total number of instances (Table 4 in Appendix G).
@@ 4.5.2 Human Alignment Evaluation
To obtain the rankings, we first calculate LFR for human annotators and for each judge-generator model pair on each set of counterfactuals generated by given generator models, as reported in Table 4 of Appendix G. Next, we compute the difference ($\Delta$) between the human LFR and the LFR of each pair. A smaller difference indicates a better ranking (lower ranking value). This process results in a ranking for each judge-generator pair. Since these pairs correspond to specific relationships $\mathcal{R}(\mathcal{G},\mathcal{J})$, we average the rankings of all pairs sharing the same relationship to obtain the average ranking per relationship, as shown in Table 3 of Appendix G. Finally, we average these rankings for each generator model to produce the overall average rankings presented in Figure 2.
@@ 5 Results
As shown in Figure 3(a) and Figure 3(b), $LLM_{\mathcal{J}}$ with $\mathcal{R}_{imwo}$ relationship achieves the highest alignment with human annotators ($rank=4.15$, $r_{m/l}=3.5$, $\rho=0.47$). In contrast, $\mathcal{R}_{dm}$ ($rank=5.58$, $r_{m/l}=1$, $\rho=0.38$) or $\mathcal{R}_{imw}$ ($rank=5.86$, $r_{m/l}=0.23$, $\rho=0.29$) demonstrate poor alignment with human judgments. This can be attributed to data contamination Li et al. (2025), as these models are either fine-tuned on the target dataset or share architectural similarities with $LLM_{\mathcal{G}}$. Such overlap could bias $LLM_{\mathcal{J}}$ in its evaluation of label flipping, potentially leading to either overestimating or underestimating the LFR. Notably, ensembling results from all judge models does not necessarily lead to better alignment, as it partially relies on results from suboptimal judge models. Additionally, we find that the first two measures (average ranking and $r_{m/\ell}$) are moderately correlated, with a Spearman correlation coefficient of 0.67, indicating that $r_{m/\ell}$ also serves as a reliable metric to capture the alignment between model pairs and human judgments.
@@
We investigate how the choice of relationship affects CDA outcomes (Appendix H). We notice that when the LFR of $LLM_{\mathcal{J}}$, due to its relationship to $LLM_{\mathcal{G}}$, aligns more closely with user study outcomes, the labels identified by $LLM_{\mathcal{J}}$ are associated with improved performance and greater robustness on both unseen and out-of-distribution data Kaushik et al. (2020); Gardner et al. (2020) (Appendix H). This association is particularly evident when these identified labels are treated as the ground-truth labels for counterfactuals, which subsequently serve as data points for CDA. This can be attributed to the fact that noisy and incorrect labels provided by judge models with suboptimal relationships contribute to performance deterioration Zhu et al. (2022); Song et al. (2023).
@@ 6 Conclusion
In this work, we emphasize the importance of the relationship between the counterfactual generator model and label flipping judge model in achieving LFR that align more closely with human annotations and in improving CDA outcomes. We further demonstrate that high downstream performance does not necessarily imply an effective judge model. Through extensive experiments, we identify that label flipping remains highly challenging across all selected tasks. Additionally, the gap between the optimal relationship and the user study is considerably large, which indicates full automation of CDA falls short and human intervention should be considered.
@@ G.3 A Sanity Check
In our user study, we randomly select 45 counterfactuals from each dataset, generated individually by five counterfactual generator models $LLM_{\mathcal{G}}$4.4). To perform a sanity check and validate the representativeness of these subsets, we conduct an additional automatic evaluation on the subset of input data from which the 45 counterfactuals were generated.